OpenStudy (anonymous):

The slope of the line below is -1. Write the point slope equation of the line using the coordinates of the labeled point (7,-7) heres what I got Y=-1x+0 is this correct ?

zepdrix (zepdrix):

You have the correct equation in `slope-intercept form`. They want it `point-slope form` though :)\[\large\rm y-y_1=m(x-x_1)\]And we were given that: \(\large\rm (x_1,y_1)=(7,-7)\) and \(\large\rm m=-1\)

OpenStudy (anonymous):

oooh whoops im not the best in math lol

zepdrix (zepdrix):

Err maybe it makes more sense to write our point-slope form in this way,\[\large\rm m(x-x_1)=y-y_1\]easier to plug in the stuff that way.

zepdrix (zepdrix):

That's ok :) Just keep eating cookies, you'll succeed.

zepdrix (zepdrix):

\[\large\rm (\color{purple}{x_1},\color{royalblue}{y_1})=(\color{purple}{7},\color{royalblue}{-7})\]\[\large\rm \color{orangered}{m=-1}\]So just plug everything in, understand?\[\large\rm \color{orangered}{m}(x-\color{purple}{x_1})=y-\color{royalblue}{y_1}\]

OpenStudy (anonymous):

so I got y=-1x thats not right is it ?